Table 4 illustrates the quantities, marginal costs, average variable costs, and average costs of a competitive firm. Refer to table 4. How much is the unit profit at price =$30?  [the problem is based on concepts in Module 5].


Table 4

Quantity (Q)

Marginal costs (MC)

Average variable costs (AVC)

Average costs (AC)

0

1

28.00

28.00

$68

2

20.00

24.00

$44

3

16.00

21.33

$35

4

14.00

19.50

$30

5

12.00

18.00

$26

6

17.83

17.83

$25

7

24.00

18.43

$24

8

30.00

19.88

$25

9

40.00

22.11

$27

10

54.00

25.30

$29

0

$5

$10

$15
Answer #1

$5

Explanation: The firm maximizes profit when price =MC. Here, price is $30. So, price = MC happens when 8 units are produced.

When 8 units are produced, total revenue = price * quantity = $30 * 8 = $240

When 8 units are produced, total cost =AC * quantity = $25 * 8 = $200

So, profit = $240 - $200 = $40.

So, profit per unit = profit/quantity = $40/8 = $5.